How many liters of 15% acid solution must be added to a 120 liters of 35% acid solution to obtain a solution that is 22%. (use Fraction for answer)
Answer by mananth(16946)

percent ---------------- quantity
Acid solution I 35 ---------------- 120 liters
Acid solution II 15 ---------------- x liters
Mixture 22 ---------------- 120 + x liters

35 * 120 + 15 x = 22 ( 120 + x )
4200 + 15 x = 2640 + 22 x
15 x + -22 x = 2640 - -4200
-7 x = -1560
/ -7
x = 222 6/7
222 6/7 liters of 15 % Acid solution II